OpenStudy (anonymous):

Joe is 10 years older than James. In 8 years twice Joe's age will equal three times James' age. How old is each now?

OpenStudy (anonymous):

their ages right now James: x Joe: x+10 In eight years: 2(x+10)=3x 2x+20=3x 20=x In eight years, James will be 20, and Joe will be 30. So now James is 12 and Joe is 22.

OpenStudy (lilg132):

James = x Joe = x + 10 8 years time: 2(x + 10) = 3x 2x + 20 = 3x 20 = x James = 20 Years old This is in 8 years time Now in the present James = 20 - 8 James = 12 Years old Joe = 12 + 10 Joe = 22 Years old Joe in 8 years = 22 + 8 * 2 Joe = 30 * 2 Joe = 60 This is in 8 years time should equal to 3 times James's age James 12 + 8 = 20 * 3 = 60 So it is correct

OpenStudy (anonymous):

thanks!

OpenStudy (lilg132):

your welcome